# Patterns of Non-Stationary Solutions to Symmetric Systems of Second-Order Differential Equations

### v2

@@ 3.1 The $s$-Folding Homomorphism
$=(G)-y_{m,j_{1}}({}^{s_{1}}H)-y_{m,j_{2}}({}^{s_{2}}H)$
$\quad+y_{m,j_{1}}y_{m,j_{2}}|W(H)|({}^{\operatorname{gcd}(s_{1},s_{2})}H)[\operatorname{gcd}(s_{1},s_{2})\in\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)]+a$ (26)
where $a\in A(G)$ satisfies $\operatorname{coeff}^{{}^{s}H}(a)=0$ for all $s\in\mathbb{N}$. In order to demonstrate each of the boolean conditions, we consider the following cases:
@@ 1. $(i)$
    Supposing that $s_{0}\notin\{s_{1},s_{2}\}$, if it is also the case that both $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are odd and $\operatorname{gcd}(s_{1},s_{2})=s_{0}$, and $s_{0}\in\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)$ then one has
@@
    and the result follows from the fact that $x_{0}|W(H)|=2$. On the other hand, if either of $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are even or if $\operatorname{gcd}(s_{1},s_{2})\neq s_{0}$, or if $s_{0}\notin\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)$ then the result follows immediately from the assumption that $s_{0}\notin\{s_{1},s_{2}\}$.
@@ 2. $(ii)$
    Supposing instead that $s_{0}\in\{s_{1},s_{2}\}$, if it is also the case that $\dim\mathcal{V}_{m,j_{0}}^{H}$ is even, then one has $y_{m,j_{0}}=0$ and the result follows from (5). If one has $s_{0}\in\{s_{0},s_{1}\}$, with both $\dim\mathcal{V}_{m,j_{1}}^{H}$, $\dim\mathcal{V}_{m,j_{s_{2}}}^{H}$ odd and $s_{0}=\gcd(s_{1},s_{2})$ and $s_{0}\in\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)$, then (5) becomes
@@
    where $a^{\prime}\in A(G)$ is such that $\operatorname{coeff}^{{}^{sm}H}(a^{\prime})=0$, and the result follows. On the other hand, if either of $\dim\mathcal{V}_{s_{1}m,j_{s_{1}}}^{H},\dim\mathcal{V}_{s_{1}m,j_{s_{1}}}^{H}$ are even (in which case one has $y_{m,j_{1}}y_{m,j_{2}}=0$) or if $\operatorname{gcd}(t,l)\neq s_{0}$, or if $s_{0}\notin\mathfrak{N}_{s_{1}}(H)\cap\mathfrak{N}_{s_{2}}(H)$, then the result follows from (5).
@@
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right],$
@@
$=\sum\limits_{I\subset\{s_{1},s_{2},\ldots,s_{N}\}}|W(H)|^{|I|-1}\left(\prod\limits_{s\in I}-y_{m,j_{s}}\right)\left[\gcd(I)\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]({}^{\operatorname{gcd}(I)}H),$
@@
$n_{s_{0}}=\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset\end{subarray}}|W(H)|^{|I|-1}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right][s_{0}=\operatorname{gcd}(I)]\prod\limits_{s\in I}-y_{m,s}.$ (29)
@@
$n_{s_{0}}$ $=\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset\end{subarray}}|W(H)|^{|I|-1}(-x_{0})^{|I|}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right]$
@@
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right],$
@@ 6 On the Non-Triviality of the Degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset S(H)\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[s_{0}\in\bigcap_{s\in I}\mathfrak{N}_{s}(H)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\mathfrak{n}^{s}(H)\right].$
@@
such the degree computation (11) becomes
@@
$\begin{cases}\ddot{x}(t)=|u|^{q}u-(L+\operatorname{Id})u,\quad t\in\mathbb{R},\;x(t)\in V\\ x(t)=x(t+p),\;\dot{x}(t)=\dot{x}(t+p),\end{cases}$ (32)
where $|u|^{q}u:=(|u_{1}|^{q}u,|u_{2}|^{q}u,\ldots,|u_{N}|^{q}u)$ for any even $q>1$ and $L:V\rightarrow V$ is the graph Laplacian matrix for an undirected graph $\mathbb{G}$ invariant under the permutation action of $D_{N}$, with $N$ vertices representing a collection of the same number of oscillating pendula and with edges representing the coupling relations between vertices. Since the natural permutation representation of $\Gamma$ on $V$
@@ 7.1 The Special Case of $\Gamma=D_{8}$
1# A G.A.P. Program for the computation of Maximal Orbit Types and Basic Degrees associated with the G-isotypic decomposition V = V_0 $\times$ V_1 $\times$ V_2 $\times$ V_3 $\times$ V_4
2LoadPackage ("EquiDeg");
3# create groups O(2)xD8xZ2
4o2:=OrthogonalGroupOverReal(2);
5d8:=pDihedralGroup(8);
6z2:=pCyclicGroup(2);
7# generate D8xZ2
8g1:=DirectProduct(d8,z2);
9# set names for subgroup conjugacy classes in D8xZ2
10SetCCSsAbbrv(g1,["Z1","Z2","D1t","D1z","D1","Z1m","Z1p","D1zt","D1pt","D1p","D2","Z4","D2t","D2zt","D2d","Z4d","D2dt","D2z","Z2p","Z4p","D4dt","D2p","D4","D2pt","D4z","D4zt","Z8","Z8d","D4t","D4d","D4p","Z8p","D8","D4pt","D8d","D8z","D8dt","D8p"]);
11# generate O(2)xD8xZ2
12G:=DirectProduct(o2,g1);
13ccs:=ConjugacyClassesSubgroups(G);
14# find Maximal Orbit Types and Basic Degrees in first O(2)-isotypic component
15irrs := Irr(G);
16# The D_8-isotypic components 0,1,2,3,4 are indexed in GAP as 1,9,11,13,7
17for i in [1,9,11,13,7] do
@@ Appendix A The $G$-Equivariant Leray-Schauder Degree
Computation of the $G$-equivariant Leray-Schauder degree. Let’s denote by $\{\mathcal{U}_{i}\}_{i\in\mathbb{N}}$ the set of all irreducible $G$-representations and define the $i$-th basic degree as follows
@@
can be practically computed, using the recurrence formula (38), as follows

### v3

@@ 3.1 The $s$-Folding Homomorphism
$=(G)-y_{m,j_{1}}({}^{s_{1}}H)-y_{m,j_{2}}({}^{s_{2}}H)$ (27)
$\quad\quad\quad+y_{m,j_{1}}y_{m,j_{2}}|W(H)|({}^{d}H)[\mathcal{B}_{\mathfrak{m}(H)}(\{d,s_{1},s_{2}\})]+a$
where $d:=\operatorname{gcd}(s_{1},s_{2})$ and $a\in A(G)$ satisfies $\operatorname{coeff}^{{}^{s}H}(a)=0$ for all $s\in\mathbb{N}$. In order to demonstrate each of the boolean conditions, we consider the following cases:
@@ 1. $(i)$
    Supposing that $s_{0}\notin\{s_{1},s_{2}\}$, if it is also the case that both $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are odd and $\operatorname{gcd}(s_{1},s_{2})=s_{0}$, and $\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})$ then one has
@@
    and the result follows from the fact that $x_{0}|W(H)|=2$. On the other hand, if either of $\dim\mathcal{V}_{m,j_{1}}^{H}$ and $\dim\mathcal{V}_{m,j_{2}}^{H}$ are even or if $\operatorname{gcd}(s_{1},s_{2})\neq s_{0}$, or if $\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})$ does not hold, then the result follows immediately from the assumption that $s_{0}\notin\{s_{1},s_{2}\}$.
@@ 2. $(ii)$
    Supposing instead that $s_{0}\in\{s_{1},s_{2}\}$, if it is also the case that $\dim\mathcal{V}_{m,j_{0}}^{H}$ is even, then one has $y_{m,j_{0}}=0$ and the result follows from (5). If one has $s_{0}\in\{s_{0},s_{1}\}$, with both $\dim\mathcal{V}_{m,j_{1}}^{H}$, $\dim\mathcal{V}_{m,j_{s_{2}}}^{H}$ odd and $s_{0}=\gcd(s_{1},s_{2})$ and $\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})$, then (5) becomes
@@
    where $a^{\prime}\in A(G)$ is such that $\operatorname{coeff}^{{}^{sm}H}(a^{\prime})=0$, and the result follows. On the other hand, if either of $\dim\mathcal{V}_{s_{1}m,j_{s_{1}}}^{H},\dim\mathcal{V}_{s_{1}m,j_{s_{1}}}^{H}$ are even (in which case one has $y_{m,j_{1}}y_{m,j_{2}}=0$) or if $\operatorname{gcd}(t,l)\neq s_{0}$, or if $\mathcal{B}_{\mathfrak{m}(H)}(\{s_{0},s_{1},s_{2}\})$ does not hold, then the result follows from (5).
@@
$\quad\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right],$
@@
$=\sum\limits_{I\subset\{s_{1},s_{2},\ldots,s_{N}\}}|W(H)|^{|I|-1}\left(\prod\limits_{s\in I}-y_{m,j_{s}}\right)\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]({}^{\operatorname{gcd}(I)}H),$
@@
$n_{s_{0}}=\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset\end{subarray}}|W(H)|^{|I|-1}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right][s_{0}=\operatorname{gcd}(I)]\prod\limits_{s\in I}-y_{m,s}.$ (30)
@@
$\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset\end{subarray}}|W(H)|^{|I|-1}(-x_{0})^{|I|}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right]$
@@
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset\{s_{1},s_{2},\ldots,s_{N}\}\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\dim\mathcal{V}_{m,j_{s}}^{H}\right],$
@@ 6 On the Non-Triviality of the Degree $G\text{\rm-deg}(\mathscr{A},B(\mathscr{H}))$
$\quad+2x_{0}\sum\limits_{\begin{subarray}{c}I\subset S(H)\\ I\neq\emptyset,\{s_{0}\}\end{subarray}}(-2)^{|I|-2}\left[\mathcal{B}_{\mathfrak{m}(H)}(I)\right]\left[s_{0}=\operatorname{gcd}(I)\right]\left[2\nmid\prod_{s\in I}\mathfrak{n}^{s}(H)\right].$
@@
such that the degree computation (11) becomes
@@
$\begin{cases}\ddot{u}(t)=|u|^{q}u-(L+\operatorname{Id})u,\quad t\in\mathbb{R},\;u(t)\in V\\ u(t)=u(t+p),\;\dot{u}(t)=\dot{u}(t+p),\end{cases}$ (33)
where $|u|^{q}u:=(|u|^{q}u_{1},|u|^{q}u_{2},\ldots,|u|^{q}u_{N})$ for any even $q>1$ and $L:V\rightarrow V$ is the graph Laplacian matrix for an undirected graph $\mathbb{G}$ invariant under the permutation action of $D_{N}$, with $N$ vertices representing a collection of the same number of oscillating pendula and with edges representing the coupling relations between vertices. Since the natural permutation representation of $\Gamma$ on $V$
@@ 7.1 The Special Case of $\Gamma=D_{8}$
1# A G.A.P. Program for the computation of Maximal Orbit Types and Basic Degrees associated with the G-isotypic decomposition
2# V = V_0 $\times$ V_1 $\times$ V_2 $\times$ V_3 $\times$ V_4
3LoadPackage ("EquiDeg");
4# create groups O(2)xD8xZ2
5o2:=OrthogonalGroupOverReal(2);
6d8:=pDihedralGroup(8);
7z2:=pCyclicGroup(2);
8# generate D8xZ2
9g1:=DirectProduct(d8,z2);
10# set names for subgroup conjugacy classes in D8xZ2
11SetCCSsAbbrv(g1,["Z1","Z2","D1t","D1z","D1","Z1m","Z1p","D1zt","D1pt","D1p","D2","Z4","D2t","D2zt","D2d","Z4d","D2dt","D2z","Z2p","Z4p","D4dt","D2p","D4","D2pt","D4z","D4zt","Z8","Z8d","D4t","D4d","D4p","Z8p","D8","D4pt","D8d","D8z","D8dt","D8p"]);
12# generate O(2)xD8xZ2
13G:=DirectProduct(o2,g1);
14ccs:=ConjugacyClassesSubgroups(G);
15# find Maximal Orbit Types and Basic Degrees in first O(2)-isotypic component (these characterize the maximal orbit types in all O(2)-isotypic components via s-folding)
16irrs := Irr(G);
17# The D_8-isotypic components 0,1,2,3,4 are indexed in GAP as 1,9,11,13,7.
18#Looping through these indices we obtain the desired data as follows:
19for i in [1,9,11,13,7] do
[…]
@@ Appendix A The $G$-Equivariant Leray-Schauder Degree
Computation of the $G$-equivariant Leray-Schauder degree. Let’s denote by $\{\mathcal{U}_{i}\}_{i\in\mathbb{N}}$ the set of all irreducible $G$-representations and define the $i$-th basic degree as follows:
@@
can be practically computed, using the recurrence formula (39), as follows: